# BarcodeBERT: Transformers for Biodiversity Analyses

### v2

@@ D.3 Pretraining
BarcodeBERT was pretrained for 35 epochs using the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $2\times 10^{-4}$, a batch size of 128, and a OneCycle learning rate scheduler (Smith & Topin, 2017). The pretraining process utilized four NVIDIA V100 GPUs and required approximately 36 hours to complete for each experiment executed. To examine the impact of pretraining, we also trained a model from scratch on the training subset of the Seen partition without any pretraining.
D.4 Fine-tuning
@@
All baseline models and BarcodeBERT were fine-tuned for 35 epochs on the supervised training subset of the Seen partition. For larger models, a batch size of 32 was used, while smaller models (CNN, HyenaDNA and BarcodeBERT) were trained with a batch size of 128. The AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $1\times 10^{-4}$ was employed, coupled with the OneCycle learning rate scheduler (Smith & Topin, 2017). We used a single NVIDIA V100 GPU for the fine-tuning process, completing within 18 hours for most models. Notably, HyenaDNA required significantly less time to fine-tune compared to other models, as a consequence of its lightweight architecture.
D.5 Linear probe training
@@
A linear classifier is applied to the embeddings generated by all the pretrained models for species-level classification. The models’ parameters are learned using stochastic gradient descent with a constant learning rate of 0.01, momentum $\mu=0.9$, and weight $\lambda=1\times 10^{-5}$.
D.6 Zero-shot clustering
@@
We evaluated the models’ ability to group sequences without supervision using the framework from Lowe et al. (2024). Embeddings were extracted from the pretrained encoders and reduced to 50 dimensions using UMAP (McInnes et al., 2018) to enhance computational efficiency while preserving data structure. These reduced embeddings were clustered with Agglomerative Clustering (L2 distance, Ward’s linkage), using the number of true species as the target number of clusters. Clustering performance was assessed with adjusted mutual information (AMI) to measure alignment with ground-truth labels.

### v3

@@ D.4 Pretraining
BarcodeBERT was pretrained for 35 epochs using the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $\alpha=2\times 10^{-4}$, a batch size of 128, and a OneCycle learning rate scheduler (Smith & Topin, 2017). The pretraining process utilized four NVIDIA V100 GPUs and required approximately 36 hours to complete for each experiment executed. To examine the impact of pretraining, we also trained a model from scratch on the training subset of the Seen partition without any pretraining.
D.5 Fine-tuning
@@
All baseline models and BarcodeBERT were fine-tuned for 35 epochs on the supervised training subset of the Seen partition. For larger models, a batch size of 32 was used, while smaller models (CNN, HyenaDNA and BarcodeBERT) were trained with a batch size of 128. The AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $\alpha=1\times 10^{-4}$ was employed, coupled with the OneCycle learning rate scheduler (Smith & Topin, 2017). We used a single NVIDIA V100 GPU for the fine-tuning process, completed within 18 hours for most models. Notably, HyenaDNA required significantly less time to fine-tune compared to other models, due to its lightweight architecture.
To train the fully-supervised BarcodeBERT model, which ablates the pretraining stage, we trained a randomly initialized model and then followed the same training procedure as our fine-tuning process.
D.6 Linear probe training
@@
A linear classifier is applied to the embeddings generated by all the pretrained models for species-level classification. The models’ parameters are learned using stochastic gradient descent with a constant learning rate of 1, momentum $\mu\!=\!0.95$, and weight decay $\lambda\!=\!1\times 10^{-9}$.
D.7 Zero-shot clustering
@@
We evaluated the models’ ability to group sequences without supervision using a modified version of the framework from Lowe et al. (2024). Embeddings were extracted from the pretrained encoders and reduced to 50 dimensions using UMAP (McInnes et al., 2018) and cosine similarity to enhance computational efficiency while preserving data structure. These reduced embeddings were clustered with Agglomerative Clustering (L2 distance, Ward’s linkage), using the number of true species as the target number of clusters. Clustering performance was assessed with adjusted mutual information (AMI) to measure alignment with ground-truth labels.
D.8 Additional classification metrics
To complement our species- and genus-level accuracy analyses, we also report the weighted F1-score, which combines precision and recall together to create a single metric, computed (on our imbalanced barcode dataset) with sklearn’s average="weighted" setting. The baselines include alignment-based BLAST, a non-SSL CNN encoder, and off-the-shelf DNA foundation models pretrained on generic genomic datasets. These are compared against BarcodeBERT (4-4-4), our model pretrained specifically on DNA barcode data.
Table 13: Weighted F1-scores of different classification models under different evaluation strategies. Baselines are grouped into an alignment-based method (BLAST), a non-SSL CNN encoder, and generic DNA foundation models pretrained off-the-shelf. These are compared against BarcodeBERT (configured with $k\!=\!4$, 4 heads, 4 layers). Boldface indicates the best result per column, and underlines indicate second place.
Model Species-level F1 (%) (seen species) Species-level F1 (%) (linear probe) Genus-level F1 (%) (1-NN probe)
BLAST 99.66 N/A 81.77
CNN encoder 98.54 N/A 49.12
DNABERT ($k\!=\!6$) 99.53 98.48 45.09
DNABERT-2 99.70 95.61 21.59
DNABERT-S 99.74 96.85 48.00
HyenaDNA-tiny (d256) 99.12 96.21 47.98
Nucleotide Transformer 99.51 96.23 37.93
BarcodeBERT (4-4-4) 99.74 99.34 76.74
D.9 Additional performance metrics
Inference-time resource consumption for each model is summarized in Table 14. CPU consumption is broadly consistent across all the foundation models, with HyenaDNA using around 3% fewer resources than the other models. This state-space, attention-free model replaces quadratic self-attention with linear recurrence kernels, behaving like a CNN during inference and yielding minimal VRAM ($\sim$130 MB) and host RAM (4.9 GB) consumption. In contrast, models in the DNABERT family require over 11 GB of system memory and 2.3–3.0 GB of GPU memory to store large attention matrices. Both the Nucleotide Transformer and BarcodeBERT moderate those costs through different design choices. NT uses the evolutionary-scale modelling architecture (7.6 GB RAM, 1.9 GB VRAM), and BarcodeBERT uses a compact BERT encoder (6.5 GB RAM, 0.7 GB VRAM). BLAST remains the most efficient option, with zero VRAM usage and only 222 MB of host RAM, reflecting its disk-backed, optimized indexing engine. Future work may include model compression (e.g. pruning, quantization) and optimized inference kernels to further narrow the efficiency gap among deep learning approaches.
Model CPU utilization (%) RAM (MB) VRAM (MB)
BLAST 7.26 222 0
CNN encoder 22.11 5,332 98
DNABERT ($k\!=\!6$) 24.91 11,345 2,285
DNABERT-2 25.01 12,669 3,010
DNABERT-S 25.01 12,696 3,009
HyenaDNA-tiny (d256) 22.18 4,851 131
Nucleotide Transformer 24.96 7,607 1,885
BarcodeBERT 25.42 6,493 694
Table 14: Resource usage comparison of DNA models. CPU usage is normalized to a 4-core system.